# A Data-to-Product Multimodal Conceptual Framework to Achieve Automated Software Evolution for Context-rich Intelligent Applications

### v1

@@ Abstract
While AI is extensively transforming Software Engineering (SE) fields, SE is still in need of a framework to overall consider all phases to facilitate Automated Software Evolution (ASEv), particularly for intelligent applications that are context-rich, instead of conquering each division independently. Its complexity comes from the intricacy of the intelligent applications, the heterogeneity of the data sources, and the constant changes in the context. This study proposes a conceptual framework for achieving automated software evolution, emphasizing the importance of multimodality learning. A Selective Sequential Scope Model (3S) model is developed based on the conceptual framework, and it can be used to categorize existing and future research when it covers different SE phases and multimodal learning tasks. This research is a preliminary step toward the blueprint of a higher-level ASEv. The proposed conceptual framework can act as a practical guideline for practitioners to prepare themselves for diving into this area. Although the study is about intelligent applications, the framework and analysis methods may be adapted for other types of software as AI brings more intelligence into their life cycles.
@@ 1 Introduction
With the advancements of Artifical Intelligence (AI) and Natural Language Processing (NLP) in the past decades, especially the rising of Large Language Model (LLM) and multimodality learning, softwrare engineering fields welcome AI techniques to be employed to every aspects of software cycles. Meanwhile, the research of intelligent applications has continuously been a hotspot [53] because of the increasing amount of data of multimodalities generated in various domains. This type of software is designed to adapt to constantly changing scenarios of rich context [53, 51], and some examples are listed in part C of figure 1. One primary characteristic of those applications is that a great portion of their system behaviors is learned from continuous interaction with the users and environment involving detection and analysis of states and activities [44, 49, 5], unlike applications of banking or insurance with more matured and stable business logic.
@@
For the purpose of introducing HLA or intelligence to software engineering, much correlated research has been applying Machine Learning (ML) and NLP techniques in requirement analysis [6] and bug detection [10] [42], verification automation [11] [29], model-driven development [47]. ASEv methodologies expect to build an automated procedure to make changes to an application or create a new application based on old systems statuses and available data if the current version of the system cannot fulfill the requirement of the present context.
The essence of AI is fundamentally built upon data-based mathematical analysis and statistical learning. The data sometimes was referred to as the context of things. The things, such as applications, objects (e.g., camera, elevator), people, and even logic, include anything that has meaningful data that can benefit AI computing in making simple or complex context-aware decisions or recommendations, which can benefit from multimodality learning since the data are from various sources and may have different formats. Thus, ASEv for CRIA requires AI to thoroughly consider the context from both the software development side and the application running side. After the integration of various AI procedures, the final target is to produce some products that can be used by existing systems, methodologies, or human users. In software engineering areas, it can start by generating simple products like new code snippets.
@@
However, the HLA of automatic generation of new intelligent applications or features for CRIA requires researchers to have an overview of the combination between the ASEv process and the multimodality learning instead of only conquering partial of the task. A data-to-product multimodal conceptual framework proposed in this research, as a synthesis of interrelated components and variables, is designed as a response to this challenge, which can shine a light on related research. Intelligent applications are context(data)-rich, which is not only from environments but also from systems and software engineering processes. From this perspective, this framework is towards both "intelligent software" engineering and intelligent "software engineering", and the latter covers the applications that are context-rich in their software engineering process, which may not necessarily be a CRIA.
In the subsequent sections of the article, we first describe the backgrounds and related work in Section 2. Section 3 outlines the key dimensions of the conceptual framework. In section 4, we present a 3S model for categorizing the solutions for the research of ASEv. Section 5 evaluates the framework and 3S model using a detailed activity diagram of an ASEv-supporting platform. Section 6 presents future work and discussion. Finally, section 7 provides the concluding remarks.
@@ 2 Backgrounds and Related Work
A conceptual framework is an integrated way of looking at problems [27], giving a broader understanding of a research problem, or telling a bigger map of possible relationships by joining together small individual concepts [19]. In order to better present the problem of achieving ASEv for CRIS, this section discusses the related concepts and literature in software evolution, data-to-product view, context dynamism, and multimodality.
@@ 2.1 Automated Software Evolution
Software evolution [25] refers to a continuous change from a lesser, simpler, or worse state to a higher or better state. To describe the HLA of CRIA, we use software evolution, a term derived from software evolution and Automated Software Engineering(ASE), to emphaise the updating and improving of the software. In this study, we consider the term to not only include the process of software design, development, maintainance, it also considers the deprecation of the software, and the generation of the software from nowhere (generation of software based on patterns learned from other software or from the old similar systems).
@@
Historically, Computer Aided Software Engineering (CASE) [45] is a classical methodology to assist designers and developers in software enginering tasks such as the development and integration. However, CASE tools were critisized with being mainly descriptive without offering theretical oreentation or attempting to explain factors affecting the adoption [18].
An interdisciplinary survey on challenges and state of the art in evolution of automated production systems (APS) was conducted by Birgit et al. [46]. They argued that fundamental methods like variability modeling, tracing etc., which enable software evolution, are until now limited to the software domain. For automated products, e.g., washing machines, and automated production systems, i.e., systems that produce (automated) products, those fundamental methods need to be adapted to the other disciplines and linked to well-known and well-established domain-specific methods like the design structure matrix in mechanical engineering. To develop those systems, not only software need to be considered, the system specific and system design which need to calcualte the hardware need to be considered throughout the whole process of the product development. Not like the hardware updates that take the cycle of decades, software may need to be adapted to market requirements or legal requirements.
@@
Martin [32] presented a survey on automatic software repair, which is to automatically find a solution to softwware bugs without human invervention. The article discusses behavioral repair where test suites, contracts, models, and crashing inputs are taken as oracle, and state repair, known as runtime repair or runtime recovery, with techniques such as checkpoint and restart, reconfiguration, and invariant restoration. This article also spans the research areas inlcuding software engineering, operating systems, programming lanugages and security.
A conceputal framework is presented by Nouredine et al. [14] for modeling the classification of software architecture evolution approaches, based on six explicit dimensions in respect to an architectural view-point as first class citizen. The proposed model is based on answering What, Where, When, Who, Why and How questions to provide a multiple choice to implement classification for architectural techniques.
@@ 2.2 Data-to-Product View
In this subsection, we discuss our particular perspective of the "Data-to-Product" view. It describes a phenomenon of a software process that utilizes the data to produce the final product instead of utilizing the data to only generate intermediate decisions or recommendations. Instead of data-to-feature or data-to-requirement, product means high-usability. So, it requires the AI techniques to be embedded into the ASEv life cycles, which will be mentioned in the framework and also illustrated in section 5.
One example of Data-to-Product can be found in Fig. 2; some new requirements can be categorized by evaluating the general features generated by learning from contextual data [43], and the auto selection or semi-auto selection can be made, and the selections will be sent to the decision making and artifact generation step, where can send a specific request based on combining extra considerations to form a loop of the decision-making process to improve the decision making and artifact generation. The special request can ask for a new round of learning based on different features from a specific data set, or based on the feedback from other phases of product generation, or the result of the feature integration. This involves a process of optimization using a learning loop. Of course, the structure can be different. But this structure is just a demonstration that the product will be one that is final or approaching the final version. This view requires a constant collection of contextual data from SE cycles, applications, as well as environments.
Figure 2: auto / semi-auto software engineering decision-making process
More specifically, all types of data that can be used to improve the application or the development process can be counted as related context in Fig 2. These data will be sent to an unimodal or multimodal learning model (more info can be found in Figure 3). Some initial features can be learned from the data. After another round of learning through LLM or expert system, decisions can be made about the generation of products automatically or semiautomatically if human experts are needed in the decision-making process. The effect from human experts can be predefined, like using a weight-based mechanism. According to the source data, different types of products can be generated, such as new requirements, new code snippets, new test cases, etc. After combining those new products, even an application can be generated using an automated way of learning from source data of requirements, code artifacts, configuration, and the relationship data between them. The power of product generation and evolvement capability is unlimited, given enough raw data.
@@ 2.3 White-Box, Black-Box, and Gray-Box Machine Learning Models
Black-box or white-box are software engineering terms describing whether the inner logic is known to the examiners or not. To learn from the context of the whole software evolution process, ML models as well as how to analyze, model, and encode the context data classification of the ML model using X-box As summarized by Michael et al. [1]: black-box ML techniques are methods that produce models which are functions of the inputs and produce outputs, where the internal functioning of the model is either hidden or too complicated to be analyzed. White-box modeling, on the contrary, produces models whose structure is not hidden but can be analyzed in detail. Decision tree models, linear regression models, bayesian networks, and fuzzy cognitive maps are the common examples of white-box models [13, 38]; in comparison, deep neural networks, support vector machines, and Large Language Models (LLM) are the common examples of black-box models [40].
@@
A most recent gray-box research is from Pan et al. [34] and presents a forward-looking roadmap for the unification of LLMs and KGs. Their roadmap consists of three general frameworks, namely, 1) KG-enhanced LLMs, which incorporate KGs during the pre-training and inference phases of LLMs or for the purpose of enhancing understanding of the knowledge learned by LLMs; 2) LLM-augmented KGs, that leverage LLMs for different KG tasks such as embedding, completion, construction, graph-to-text generation, and question answering; and 3) Synergized LLMs + KGs, in which LLMs and KGs play equal roles and work in a mutually beneficial way to enhance both LLMs and KGs for bidirectional reasoning driven by both data and knowledge.
@@ 2.4 Context Dynamism
One primary characteristic of those applications is that a great portion of their system behaviors is learned from continuous interaction with the users and environment involving detection and analysis of states and activities [44, 49, 5], unlike applications of banking or insurance with more matured and stable business logic.
The feature of those applications is that they are based on behaviors, not only on a very mature business. For matured business applications, designers need to analyze the business model and the data model and then make decisions about which data model is the best design. Some behaviors-based applications can not adapt to the new normal very well. Even for matured business applications, big changes can be rare, but some other decisions can be made, so it can focus on the improvements of the flow, the adding of a new feature, or the enhancement of an old feature. So, the main focus is on the fast-changing context and fast-changing business applications. Based on various types of data formats, various data inputs, and other things, all those things need to be adapted, and then the code needs to be changed. In order to study this, we need a conceptual framework to support the building of this automated process. So this conceptual framework will not only cover the context dynamism of a lot of applications but also cover all that is context-rich since they may share similarities, even if the matured businesses are steady.
For CRIA, contextual data can change frequently, and so can decision processes. Thus, contextual reasoning logic needs to adapt to the changes accordingly. The dynamism can sometimes lead to a new system feature [51]. For example, a smart elevator ships people in most scenarios, but for a period of time, it sees people convoying some large devices. Then, how it updates itself to fit the new normal and many more new normals could be a challenging problem. Some other types of dynamism can be perceived given enough data and time for analysis. Thus, the software team or a designated system should be smart enough to take care of updating the intelligent elevator system by automatically identifying the dynamism.
Thus, context dynamism refers to the dynamic aspects of entities, attributes, and states and the corresponding relationship changes within intelligent system domains. Below are some examples in a smart campus setting; however, additional instances could exist in real-world scenarios.
1) New Joining of Entity, Attribute, or State, e.g., a new student is joining a smart campus.
2) Relationship Changes, e.g., a student enrolls in a class provided by a professor.
3) Hot-spot Situation Forming, e.g., a student makes a decision only with the accompany of the happening of another fact.
4) Cycles and steps of transitions affecting decision-making: a student will visit a place or make a decision only after a pattern of visiting some other places.
@@ 2.5 Multimodality
Multimodality is extensively studied in AI / ML research [31, 2, 37] and industry engineering [15, 17]. It refers to a phenomenon of harnessing data from heterogeneous sources to achieve a better understanding of some target data or make a more precise decision by reasoning those data from multimodalities. Due to the intrinsic similarity, the popularity of multimodality can be derived from the data fusion [9, 23], while multimodality is a natural fact that exists in our daily life, as Tadas et. al pointed out in their article [2]: "Our experience of the world is multimodal - we see objects, hear sounds, feel the texture, smell odors, and taste flavors."
@@
A definition of multimodality for ML was also given by Letitia et al. [35]: A ML task is multimodal when inputs or outputs are represented differently or are composed of distinct types of atomic units of information. In this perspective, to achieve automated software evolution by utilizing data with different formats from various phases of the software engineering is a multimodal learning process, as shown in Figure 3, which contains a structure adapted from the work [36]. We can give a definition of multimodality in software engineering as the term to describe the learning process from heterogeneous sources of data from various phases of software engineering, the environment, the internal environment of applications, and all related models and text data of various formats.
@@
Figure 4 demonstrates a Data-to-Product multimodal framework for ASEv. It includes dimensions of Context Sources (C), Data Modalities (D), Multimodal Learning (M), Key features of ASEv (K), and Products (P). The internal relations among them are as follows: C provides data to D, D is the source of M, K is the key feature of ASEv, and M. P is the result of applying K. However, the process of generating P is not trivial and needs a tailored platform to achieve the task. In section 5, we introduce an activity diagram of such type of platform to achieve this connection.
The items listed in the frameworks dimensions are not exhaustive, and they are examples that should be considered with higher priorities to the author’s best knowledge. In the following subsections, we analyze key dimensions, some exemplified items, and features of the framework.

### v5

# A Multimodal Conceptual Framework to Achieve Automated Software Evolution for Context-rich Intelligent Applications

@@ Abstract
While AI is extensively transforming Software Engineering (SE) fields, SE is still in need of a framework to consider overall all phases to facilitate Automated Software Evolution (ASEv), particularly for intelligent applications that are context-rich instead of conquering each division independently. Its complexity comes from the intricacy of the intelligent applications, the heterogeneity of the data sources, and the constant changes in the context. This study proposes a conceptual framework for achieving automated software evolution, emphasizing the importance of multimodality learning. A Selective Sequential Scope Model (3S) model is developed based on the conceptual framework, and it can be used to categorize existing and future research when it covers different SE phases and multimodal learning tasks. This research is a preliminary step toward the blueprint of a higher-level ASEv. The proposed conceptual framework can act as a practical guideline for practitioners to prepare themselves for diving into this area. Although the study is about intelligent applications, the framework and analysis methods may be adapted for other types of software as AI brings more intelligence into their life cycles.
@@ 1 Introduction
With the advancements of Artificial Intelligence (AI) and Natural Language Processing (NLP) in the past decades, especially the rise of the Large Language Model (LLM) and multimodality learning, software engineering fields welcome AI techniques to be employed in every aspect of software cycles. Meanwhile, the research of intelligent applications has continuously been a hotspot [73] because of the increasing amount of data of multimodalities generated in various domains. This type of software is designed to adapt to constantly changing scenarios of rich context [73, 70], and some examples are listed in part C of figure 1. One primary characteristic of those applications is that a significant portion of their system behaviors is learned from continuous interaction with the users and environment involving detection and analysis of states and activities [62, 67, 10], unlike applications of banking or insurance with more matured and stable business logic.
@@
For the purpose of introducing HLA or intelligence to software engineering, lots of correlated research has been applying Machine Learning (ML) and NLP techniques in requirement analysis [12] and bug detection [17] [56], verification automation [18] [42], model-driven development [65]. ASEv methodologies expect to build an automated procedure to make changes to an application or create a new application based on old systems statuses and available data if the current version of the system cannot fulfill the requirement of the present context.
One important essence of AI is fundamentally built upon data-driven mathematical analysis, logical reasoning, statistical learning, and algorithmic search. The data sometimes was referred to as the context of things. The things, such as applications, objects (e.g., camera, elevator), people, and even logic, include anything that has meaningful data that can benefit AI computing in making simple or complex context-aware decisions or recommendations, which can benefit from multimodality learning since the data are from various sources and may have different formats. Thus, ASEv for CRIA requires AI to thoroughly consider the context from both the software development side and the application running side. After the integration of various AI procedures, the final target is to produce some products that can be used by existing systems, methodologies, or human users. In software engineering areas, it can start by generating simple products like new code snippets.
@@
However, the HLA of generation of new intelligent applications or features for CRIA requires researchers to have an overview of the combination between the ASEv process and the multimodality learning instead of only conquering partial of the task. A data-to-product multimodal conceptual framework proposed in this research, as a synthesis of interrelated components and variables, is designed as a response to this challenge, which can shine a light on related research. A conceptual framework is an integrated way of looking at problems [38], giving a broader understanding of a research problem, or telling a bigger map of possible relationships by joining together small individual concepts [28]. Intelligent applications are rich in contextual data, which is not only from environments but also from systems and software engineering processes. From this perspective, this framework is towards both "intelligent software" engineering and intelligent "software engineering," and the latter covers the applications that are context-rich in their software engineering process, which may not necessarily be a CRIA.
The subsequent sections of the article are organized as follows: Section 2 describes the background and related work. Section 3 outlines the key dimensions of the conceptual framework. Section 4 introduces a 3S model for categorizing solutions in the research of ASEv. Section 5 evaluates the framework and the 3S model by applying them to ASEv-related research. Section 6 discusses related topics of interest, limitations, and concerns. Finally, Section 7 provides the concluding remarks.
@@ 2 Backgrounds and Related Work
In order to better present the problem of achieving ASEv for CRIS, this section discusses the related concepts and literature in software evolution, data-to-product view, context dynamism, and multimodality.
@@ 2.1 Automated Software Evolution
Software evolution refers to "a continuous change from a lesser, simpler, or worse state to a higher or better state" [36]. To describe the HLA of CRIA, automated software evolution, a term derived from software evolution and Automated Software Engineering(ASE), is used to emphasize the updating and improving of the software. In this study, the term does not only include the process of software design, development, and maintenance; it also considers the deprecation of the software and the generation of the software from nowhere (generation of software based on patterns learned from other software or from the old, similar systems).
@@
Historically, Computer Aided Software Engineering (CASE) [63] is a traditional approach that aids designers and developers in software engineering activities, including development and integration. However, CASE tools were criticized for being mainly descriptive. They provide neither a theoretical framework nor an attempt to explain the factors influencing adoption [27].
An interdisciplinary survey on challenges and state of the art in the evolution of automated production systems (APS) was conducted by Birgit et al. [64]. They argued that fundamental techniques such as variability modeling and tracing, which facilitate software evolution, are confined to the software domain. To develop those APS, not only software needs to be considered, but the system specifics and system design, which need to calculate the hardware, need to be considered throughout the whole process of product development. Although their research targets are more related to mechatronic automated products, such as washing machines or other automated systems that produce automated products, the perceived underlying rule can be applied to other types of software and hardware pairs. Take the iPhone’s hardware ecosystem and software ecosystem as an example; not only do the operating system updates happen after hardware updates, but those mobile applications will also have to adapt themselves to the new features provided by new hardware and software. A vanward idea about hardware considers it as a new type of software [3]. Software often needs to be revised to meet changing market and regulatory demands.
@@
Martin [45] presented a survey on automatic software repair to resolve software bugs autonomously without interventions from human experts. The article discusses behavioral repair and state repair. Behavioral repair treats test suites, contracts, models, and crashing inputs as an oracle. State repair works at runtime using checkpoints and restarts, system reconfiguration, and invariant rehabilitation techniques. This survey work notably covers an extensive array of concepts in automatic software repair and spans a broad range of research areas, including not only software engineering but also operating systems, programming languages, and security.
A conceptual framework is presented by Nouredine et al. [21] for modeling the classification of software architecture evolution approaches based on six explicit dimensions (What, Where, When, Who, Why, and How questions) treating the architectural viewpoint as a primary focus. Taking "When" as an example, it refers to the time of evolution, which consists of design-time evolution and runtime evolution. The former is at the design stages of software, allowing improvement and extension of the architecture by predicting the evolution earlier. The latter, as considered a primary theme in architecture adaptation, encompasses evolving at compile-time, load time, and dynamic time. As a comparison, this research studies a conceptual framework for resolving how multimodalities and AI shape the future automatic software evolution, which encloses architectural evolution.
@@ 2.2 Data-to-Product View
This subsection discusses a particular perspective of the "Data-to-Product" view, which describes a phenomenon of a software process that utilizes the data to produce the final product instead of utilizing the data to only generate intermediate decisions or recommendations. Compared with "requirement" and "feature" in terms like "data-to-feature" and "data-to-requirement," "product" is the final result of the software evolution and means high usability. So, it requires the AI techniques to be embedded into all ASEv life cycles, which will be mentioned in the framework and also illustrated in section 5.
Figure 2: an auto / semi-auto software engineering decision-making process
One example of Data-to-Product can be found in Fig. 2; some new requirements can be categorized by evaluating the general features generated by learning from contextual data [60]. Auto selections or semi-auto selections can be made, and the selections will be sent to the decision-making and artifact generation step, where it can send a specific request based on combining extra considerations to form a loop of the decision-making process to improve the decision-making and artifact generation. The special request can ask for a new round of learning based on different features from a specific data set, feedback from other phases of product generation, or the result of the feature integration. This involves a process of optimization using a learning loop. Of course, the structure can be different. But this structure is just a demonstration that the product will be one that is final or approaching the final version. This view requires a constant collection of contextual data from SE cycles, applications, as well as environments.
More specifically, all types of data that can be used to improve the application or the development process can be counted as related context in Fig 2. These data will be sent to an unimodal or multimodal learning model (more info can be found in Figure 3). Some initial features can be learned from the data. After another round of learning through LLM or expert system, decisions can be made about the generation of products automatically or semiautomatically if human experts are needed in the decision-making process. The effect from human experts can be predefined, like using a weight-based mechanism. According to the source data, different types of products can be generated, such as new requirements, new code snippets, new test cases, etc. After combining those new products, an application can even be generated using an automated way of learning from source data of requirements, code artifacts, configuration, and the relationship data between them. The power of product generation and evolvement capability is unlimited, given enough raw data.
@@ 2.3 White-Box, Black-Box, and Gray-Box Machine Learning Models
Black-box or white-box are software engineering terms describing whether the inner logic is known to the examiners or not. ML models are classified using X-Box (white, black, or gray-box) w.r.t how to analyze, model, and encode the context data. As summarized by Michael et al. [1]: black-box ML techniques refer to methods that generate models where the internal workings are either concealed or too complex to be analyzed, producing outputs based on the inputs. In contrast, white-box modeling involves models with transparency that can be analyzed in detail. Decision tree models, linear regression models, bayesian networks, and fuzzy cognitive maps are the common examples of white-box models [20, 52]; in comparison, deep neural networks, support vector machines, and Large Language Models (LLM) are the common examples of black-box models [54].
@@
A most recent gray-box research is from Pan et al. [48] and outlines a forward-looking roadmap for integrating Large Language Models (LLMs) and Knowledge Graphs (KGs). This roadmap features three primary frameworks: Firstly, KG-enhanced LLMs, which integrate KGs during both the pre-training and inference phases of LLMs to improve their understanding of acquired knowledge; Secondly, LLM-augmented KGs, which utilize LLMs to perform various KG tasks such as embedding, completion, construction, graph-to-text generation, and question answering; and thirdly, Synergized LLMs + KGs, where LLMs and KGs collaborate equally to enhance each other, facilitating bidirectional reasoning that leverages both data and knowledge.
Takeishi et al. [59] define deep gray-box models as compositions of data-driven models and theory-driven models, with the former being deep neural networks and the latter being human-understandable models with a theory as a backbone. They empirically analyze the sum of regularizers, including those that can prevent the theory-driven models from being ignored.
@@ 2.4 Context Dynamism
One primary characteristic of CRIA is that a great portion of their system behaviors is learned from continuous interaction with the users and environment involving detection and analysis of states and activities [62, 67, 10], unlike applications of banking or insurance with more matured and stable business logic.
For CRIA, contextual data can change frequently, and so can the decision processes of the applications. Thus, contextual reasoning logic needs to adapt to the changes accordingly. The dynamism can sometimes lead to a new system feature [70]. For example, a smart elevator ships people in most scenarios, but for a period of time, it sees people convoying some large devices. Then, how it updates itself to fit the new normal and many more new normals could be a challenging problem. Some other types of dynamism can be perceived given enough data and time for analysis. Thus, the software team or a designated system should be smart enough to take care of updating the intelligent elevator system by identifying the dynamism automatically.
Thus, context dynamism refers to the dynamic aspects of entities, attributes, and states and the corresponding relationship changes within intelligent system domains [69]. One type of context dynamism can be the new joining of entity, attribute, or state, and an exemplified instance is "a new student is joining a smart campus." Other types include relationship changes, hotspot situation forming, and cycles and steps of transitions affecting decision-making [69].
Although big changes for business-matured applications can be rare, some small or medium changes due to context changes are unavoidable, such as adaptation to regularities, new versions of the underlying software, or optimized workflow from analyzing contextual data. Although the main focus of ASEv in this study is CRIA, it can also be applied to business-stable applications. Changes in the data can result in changes in the code or configuration. A conceptual framework is needed to support the building of this automated application evolution process.
@@ 2.5 Multimodality
Multimodality is extensively studied in AI / ML research [44, 2, 51] and industry engineering [22, 26]. It refers to a phenomenon of harnessing data from heterogeneous sources to achieve a better understanding of some target data or make a more precise decision by reasoning those data from multimodalities. Due to the intrinsic similarity, the popularity of multimodality can be derived from the data fusion [15, 34], while multimodality is a natural fact that exists in our daily life, as Tadas et al. pointed out in their article [2]: "Our experience of the world is multimodal - we see objects, hear sounds, feel the texture, smell odors, and taste flavors."
@@
A definition of multimodality for ML was also given by Letitia et al. [49]: A ML task is multimodal when inputs or outputs are represented differently or are composed of distinct types of atomic units of information. In this perspective, to achieve automated software evolution by utilizing data with different formats from various phases of the software engineering is a multimodal learning process, as shown in Figure 3, which contains a structure adapted from the work [50].
Thus, multimodal learning in software engineering can be defined as the term to describe the learning process with data from heterogeneous sources of various phases of software engineering. Data includes text, documents, code, images, videos, models, etc., from external and internal environments of applications. The definition emphasizes the differences between sources: as long as the sources are different, they contribute to multimodality, regardless of whether their formats are the same or distinct.
@@
Figure 4 demonstrates a Data-to-Product multimodal framework for ASEv. It includes dimensions of Context Sources (C), Data Modalities (D), Multimodal Learning (M), Key features of ASEv (K), and Products (P). The internal relations among them are as follows: C provides data to D, D is the source of M, K is the key feature of ASEv, and M. P is the result of applying K. The process of generating P is not trivial and can need a tailored platform to achieve each of the tasks.
The items listed in the framework dimensions are not exhaustive, and they are examples that should be considered with higher priorities to the author’s best knowledge. In the following subsections, key dimensions, exemplified items, and features of the framework are analyzed.